OpenStudy (anonymous):

find the reference angle of the given angle: 150 degrees(more info in comments)

OpenStudy (anonymous):

here's the example from my textbook it said \[A=1/2ab \sin \theta\] =1/2(10)(3)sin 120 = 15 sin 60 15\[\sqrt{3}/2\] (they multiplied) = about 13cm^2 but then i don't understand when the first problem says find the reference angle for the given angle